# DRAGON: Dynamic RAG Benchmark On News

### v2

@@ 6 Experiments
To assess the quality of the generated question-answer pairs, a human evaluation study is conducted. Each QA pair from Sandbox Datasets (Sec. 2) is independently evaluated by 3 expert annotators along the evaluation criteria from Appx. F. Annotators were asked to mark each pair as “Good” or “Not Good” with respect to each criterion (an example of the assessment instruction is described in Appx. I). To account for potential subjectivity in judgment, we considered a QA pair to be acceptable with the majority vote.
Tab. 1 shows the proportion of QA pairs considered good for each dataset version and each evaluation criterion. The results establish the high quality of generated questions and significant context dependency. The answer evaluation proved the prevalence of correct answers, while the answer uniqueness is lower, so the ground truth answer can be substituted with another entity from the text. This fact exhibits the importance of LLM-as-Judge evaluation for RAG systems to avoid rephrasing influence.
| Criterion | Apr | May | Jun | |
| Question | Question Literacy | 0.96 | 0.97 | 0.99 |
| Clarity | 0.99 | 1.00 | 1.00 | |
| Naturalness | 0.98 | 0.96 | 0.97 | |
| Context Sufficiency | 0.98 | 0.98 | 0.99 | |
| Context Necessity | 0.95 | 0.97 | 0.98 | |
| Answer | Correctness | 0.95 | 0.92 | 0.96 |
| Uniqueness | 0.76 | 0.78 | 0.80 | |
@@
Retrieval evaluation results presented in Tab. 2 demonstrate consistently strong performance across all evaluated retriever models. Among them, Qwen3Embedding 8B and E5 Mistral7b Instruct achieve the highest scores, performing practically on par. mE5lInstruct also performs competitively. As for FRIDA, it also demonstrates strong performance but its results are slightly inferior to those of the competitors.
| Retriever | Hit Rate | Recall | NDCG |
| FRIDA | 0.88 | 0.84 | 0.85 |
| mE5 | 0.90 | 0.87 | 0.86 |
| Qwen 3 | 0.93 | 0.90 | 0.89 |
| E5 Mistral | 0.93 | 0.89 | 0.87 |
@@
Tab. 4 presents general results of the full RAG-system evaluation and detailed criteria-wise evaluation scores are given in Appx. J.
First, it can be seen that the choice of the retrieval model plays a crucial role and same as in Sec. 6. Qwen3 and E5 Mistral show the strongest results. Second, it should be noted that the general LLM ranking remains the same with every retrieval, with Qwen 2.5 and Zero Mistral heading the list by Judge Score and SM, and Gemma 3 over-performing other competitors by Rouge-L.
Overall, the results show that classic metrics such as Rouge-L are not objective enough and do not allow evaluating all aspects of the RAG task. In general, system scores positively characterize DRAGON as being complex enough for modern RAG-systems, allowing researchers to evaluate their capabilities at a high level. In the future, we also plan to complexify Judge Evaluation criteria, thus providing an opportunity for an adequate assessment of more advanced models than those that exist nowadays and avoiding the danger of the benchmark being solved.
| Retriever | LLM | Rouge-L | SM | Judge Score |
| FRIDA | Qwen 2.5 | 0.42 | 0.47 | 0.83 |
| Qwen 2.5 | 0.33 | 0.43 | 0.80 | |
| Ruadapt Qwen | 0.40 | 0.45 | 0.82 | |
| Zero Mistral | 0.44 | 0.45 | 0.83 | |
| Gemma 3 | 0.44 | 0.44 | 0.81 | |
| Gemma 3 | 0.40 | 0.47 | 0.81 | |
| Qwen 3 | Qwen 2.5 | 0.44 | 0.51 | 0.86 |
| Qwen 2.5 | 0.36 | 0.45 | 0.83 | |
| Ruadapt Qwen | 0.42 | 0.48 | 0.86 | |
| Zero Mistral | 0.45 | 0.48 | 0.86 | |
| Gemma 3 | 0.47 | 0.47 | 0.84 | |
| Gemma 3 | 0.43 | 0.49 | 0.85 | |
| E5 Mistral | Qwen 2.5 | 0.44 | 0.51 | 0.86 |
| Qwen 2.5 | 0.36 | 0.46 | 0.82 | |
| Ruadapt Qwen | 0.41 | 0.47 | 0.85 | |
| Zero Mistral | 0.45 | 0.47 | 0.85 | |
| Gemma 3 | 0.46 | 0.47 | 0.83 | |
| Gemma 3 | 0.41 | 0.48 | 0.85 | |
| mE5 | Qwen 2.5 | 0.43 | 0.49 | 0.85 |
| Qwen 2.5 | 0.34 | 0.44 | 0.82 | |
| Ruadapt Qwen | 0.39 | 0.46 | 0.83 | |
| Zero Mistral | 0.44 | 0.46 | 0.84 | |
| Gemma 3 | 0.46 | 0.46 | 0.83 | |
| Gemma 3 | 0.41 | 0.48 | 0.83 |
@@ 7 Conclusion
We presented DRAGON (Dynamic RAG Benchmark On News), the first dynamic benchmark for evaluating retrieval-augmented generation systems in Russian. DRAGON reflects real-world deployment settings by leveraging a regularly updated knowledge base and focuses on the recurrent evaluation of both retriever and generator components. Our benchmark addresses the current lack of standardized RAG evaluation tools for the Russian language. We release the benchmark, the framework comprising a question generation pipeline and evaluation scripts, and launch a public leaderboard, to support reproducible, transparent, and community-driven research. In the future, with the evolving capabilities of RAG systems, we plan to extend the benchmark by introducing new question types, refining the LLM-as-Judge criteria. In addition, we aim to open-source previous snapshots of the evolving datasets to support reproducibility and foster further community research.
@@
The proposed benchmark consists entirely of Russian language documents and questions. Although the methodology itself could be easily applied to any other language, in the current state the only one language is presented.
@@
The chosen evaluation metrics, such as ROUGE, which is essentially an n-gram precision, predominantly focus on surface-level matching. These metrics may not adequately reflect the semantic and pragmatic aspects of the generated content, and have limited correlation with human judgment Deutsch et al. (2022). LLM as Judge evaluation is designed to mitigate a semantic gap of n-gram based metrics. However, RAG benchmark requires specific criteria to catch system performance details. Building more adapted judge models can improve the quality of the assessment.
@@ Appendix C Baseline Details
| Zero Mistral | 24B | ZeroAgency/Zero Mistral-24B | – |
| Gemma 3 | 12B | google/gemma-3-12b-it | Team (2025) |
| Gemma 3 | 27B | google/gemma-3-27b-it | Team (2025) |
@@
Appendix D Question Types
We use the typology of four question types inspired by Yang et al. (2024). Below each type is described in more detail:
Simple
relations: (Morty Smith | voice | Keisuke Chiba)
question: Who voiced Morty Smith?
answer: Keisuke Chiba
Set
question: What projects has Ryan Otter composed music for?
answer: Trigger, Method
Multi-Hop
answer: China
Conditional
question: Who performed at M-bar and met with Dmitry Dibrov?
answer: Roman Miroshnichenko
Appendix E Graph correspondence filtering
The question-answer pairs were compared to the graph used for generation. Each node in the graph (entity) is assigned 2 coefficients: question presence and answer presence. It is evaluated as the scaled Levenshtein distance between the name of the entity and the closest substring from the question and answer. These values allow us to check that all entities have been mentioned correctly.
Appendix F Question-Answer Evaluation Criteria
@@
Appendix G LLM-as-Judge filtering
For fine-grained filtering of the generated question-answer pairs, we employ the LLM-as-a-Judge paradigm, leveraging a POLLUX 7B Martynov et al. (2025) model to systematically evaluate each sample along a diverse set of criteria. This approach offers a scalable and cost-effective alternative to human annotation , while still maintaining a high level of judgment quality.
Every criterion from Appx. F is transformed into a separate prompt with the specific scoring scale (0-2). For answer criteria (Literacy, Correctness, Uniqueness Based on Context), the prompt contains a news article, a question, and an answer; for other criteria, the answer is omitted. The example is classified as positive according to the particular criterion if the judge model assigns a rating of 1 or higher. This threshold was chosen to imitate the majority vote used in human evaluation.
| Criterion | Precision | Recall |
| Question Literacy | 0.96 | 0.99 |
| Question Clarity | 0.99 | 0.62 |
| Question Naturalness | 0.96 | 0.52 |
| Context Sufficiency | 0.94 | 0.71 |
| Context Necessity | 0.93 | 0.95 |
| Answer Correctness | 0.95 | 0.82 |
| Answer Uniqueness Based on Context | 0.85 | 0.78 |
| Answer Literacy | 0.91 | 0.97 |
Appendix H Experimental Setup Details
H.1 Embedding Model Prefixes
| Model | Query prefix | Text prefix |
| FRIDA | search_query: | search_document: |
H.2 LLM Prompt Template
‘‘Answer the question using the provided context.
Give me only an answer.
<context> {context} </context>
Question: {question}
Answer: ’’
H.3 Model Configuration
For serving models, we used the vLLM framework. The used model parameters are shown in the Tab. 8.
| Model | tp_size | max_context_length |
| Qwen 2.5 | 4 | 32768 |
| Qwen 2.5 | 1 | 32768 |
| Ruadapt Qwen | 4 | 32768 |
| Zero Mistral | 4 | 131072 |
| Gemma 3 | 1 | 131072 |
| Gemma 3 | 4 | 131072 |
H.4 Metrics
The performance of retrieval is measured by 3 metrics:
Hit Rate measures the proportion of queries for which relevant document appears among the top-k retrieved results.
Recall computes the ratio of relevant documents successfully retrieved out of all ground-truth relevant documents.
Normalized Discounted Cumulative Gain (NDCG) evaluates the ranking quality by assigning higher scores to relevant documents retrieved at higher ranks.
We evaluate End-to-end RAG systems with:
ROUGE-L, used to assess generation quality, is based on the longest common subsequence between the model output and reference text.
Substring Matching measures whether key segments from the reference answer are present in the generated response.
Judge Score is used to evaluate the overall answer quality, is calculated as the average of the automatic scores from Pollux 151515https://ai-forever.github.io/POLLUX/ across multiple criteria (e.g., correctness, completeness, and relevance).
Appendix I Human Evaluation Interface
@@
Appendix J LLM-as-Judge RAG Evaluation Criteria

### v3

@@ 6 Experiments
To assess the quality of the generated question-answer pairs, a human evaluation study is conducted. Each QA pair from Sandbox Datasets (Sec. 2) is independently evaluated by 3 expert annotators along the evaluation criteria from Sec. D. Annotators were asked to mark each pair as “Good” or “Not Good” with respect to each criterion. To account for potential subjectivity in judgment, we considered a QA pair to be acceptable with the majority vote. Evaluation results are provided in Tab. 4.
| Criterion | Apr | May | Jun |
| Question Literacy | 0.96 | 0.97 | 0.99 |
| Clarity | 0.99 | 1.00 | 1.00 |
| Naturalness | 0.98 | 0.96 | 0.97 |
| Context Sufficiency | 0.98 | 0.98 | 0.99 |
| Context Necessity | 0.95 | 0.97 | 0.98 |
| Correctness | 0.95 | 0.92 | 0.96 |
| Uniqueness | 0.76 | 0.78 | 0.80 |
@@
Retrieval evaluation results presented in Tab. 5 demonstrate consistently strong performance across all evaluated retriever models. Among them, Qwen3Embedding 8B and E5 Mistral7b Instruct achieve the highest scores. mE5Large Instruct also performs competitively. As for FRIDA, it also demonstrates strong performance but its results are slightly inferior to those of the competitors.
| Retriever | Hit Rate | MRR |
| FRIDA | 0.932 | 0.822 |
| Qwen 3 | 0.960 | 0.867 |
| E5 Mistral | 0.956 | 0.851 |
| mE5 | 0.949 | 0.834 |
@@
| LLM | Rouge2 | RougeL | JS |
| Retrieval: FRIDA | |||
| Gemma 3 12B it | 0.14 | 0.22 | 0.63 |
| Gemma 3 27B it | 0.14 | 0.22 | 0.64 |
| Qwen 2.5 32B | 0.09 | 0.16 | 0.62 |
| Qwen 2.5 7B | 0.08 | 0.13 | 0.57 |
| Qwen3 32B | 0.08 | 0.14 | 0.64 |
| Rudadapt Qwen 32B | 0.13 | 0.22 | 0.72 |
| Retrieval: mE5 | |||
| Gemma 3 12B it | 0.15 | 0.24 | 0.67 |
| Gemma 3 27B it | 0.15 | 0.24 | 0.67 |
| Qwen 2.5 32B | 0.10 | 0.18 | 0.66 |
| Qwen 2.5 7B | 0.09 | 0.15 | 0.63 |
| Qwen3 32B | 0.11 | 0.18 | 0.69 |
| Rudadapt Qwen 32B | 0.14 | 0.21 | 0.74 |
| Retrieval: Qwen 3 | |||
| Gemma 3 12B it | 0.16 | 0.26 | 0.71 |
| Gemma 3 27B it | 0.17 | 0.26 | 0.72 |
| Qwen 2.5 32B | 0.11 | 0.19 | 0.68 |
| Qwen 2.5 7B | 0.09 | 0.16 | 0.64 |
| Qwen3 32B | 0.11 | 0.19 | 0.71 |
| Rudadapt Qwen 32B | 0.16 | 0.25 | 0.82 |
| Retrieval: E5 Mistral | |||
| Gemma 3 12B it | 0.16 | 0.25 | 0.68 |
| Gemma 3 27B it | 0.16 | 0.25 | 0.70 |
| Qwen 2.5 32B | 0.12 | 0.19 | 0.68 |
| Qwen 2.5 7B | 0.09 | 0.16 | 0.64 |
| Qwen3 32B | 0.11 | 0.19 | 0.71 |
| Rudadapt Qwen 32B | 0.16 | 0.25 | 0.79 |
The results are provided in Table 6. Overall, the results show that classic metrics such as Rouge-L are not objective enough and do not allow evaluating all aspects of the RAG task.
First, it can be seen that the choice of the retrieval model plays a crucial role. Qwen3 and E5 Mistral show the strongest results. Second, it should be noted that the general LLM ranking remains the same with every retrieval, with Rudadapt Qwen 32B heading the list by Judge Score, and Gemma 3 outperforming other competitors by Rouge metrics.
In general, system scores positively characterize DRAGON as being complex enough for modern RAG-systems, allowing researchers to evaluate their capabilities at a high level. In the future, we also plan to complexify Judge Evaluation criteria, thus providing an opportunity for an adequate assessment of more advanced models than those that exist nowadays and avoiding the danger of the benchmark being solved.
@@ 7 Conclusion
We presented DRAGOn, a method to design RAG benchmark on any periodically updated document source, with it we created the dynamic benchmark for evaluating retrieval-augmented generation systems in Russian. DRAGOn is designed for real-world deployment settings by leveraging a regularly updated knowledge base and focusing on the recurrent evaluation of both retriever and generator components. Our methodology addresses the current lack of standardized RAG evaluation tools; thus, we created a sample benchmark for the Russian language. We release the benchmark, which comprises a question generation pipeline and evaluation scripts, and launch a public leaderboard to support reproducible, transparent, and community-driven research. In the future, with the evolving capabilities of RAG systems, we plan to extend the benchmark by introducing new question types, refining the LLM-as-Judge criteria. In addition, we aim to open-source previous snapshots of the evolving datasets to support reproducibility and foster further community research.
@@
The proposed benchmark consists entirely of Russian language documents and questions. Although the methodology itself could be easily applied to any other language, in its current state, only one language is presented.
@@
The chosen evaluation metrics, such as ROUGE, which is essentially an n-gram precision, predominantly focus on surface-level matching. These metrics may not adequately reflect the semantic and pragmatic aspects of the generated content and have limited correlation with human judgment Deutsch et al. (2022). The LLM as Judge evaluation is designed to mitigate the semantic gap of n-gram-based metrics. However, the RAG benchmark requires specific criteria to capture details of system performance. Building more adapted judge models can improve the quality of the assessment.
@@
Error Analysis
A further limitation of our current benchmark release is the lack of a systematic error analysis of model failures. While we report aggregate retrieval and generation scores, we do not yet provide a fine-grained breakdown of common failure modes (e.g., retrieval misses vs. ranking issues, incomplete evidence aggregation in multi-hop questions, hallucinations under partially relevant context). Such analysis is important both for interpreting leaderboard progress and for understanding whether improvements come from better retrieval, better grounding/faithfulness, or exploiting dataset artifacts. In future work, we plan to add structured error taxonomies and identify systematic weaknesses of evaluated RAG pipelines.
@@
Acknowledgments
We would like to express our deep appreciation to Ivan Bondarenko for his contribution to our pipeline and generous support of this work. This research partially done by A.F. is an output of a research project implemented as part of the Basic Research Program at the National Research University Higher School of Economics (HSE University).
@@ Appendix C Baseline Details
| Qwen 3 | 32B | Qwen/Qwen3-32B | Team (2025b) |
| Gemma 3 | 12B | google/gemma-3-12b-it | Team (2025a) |
| Gemma 3 | 27B | google/gemma-3-27b-it | Team (2025a) |
@@
Appendix D Question-Answer Evaluation Criteria
@@
Appendix E Human Evaluation Interface
@@
Appendix F LLM-as-Judge RAG Evaluation Criteria